### v2

# Integrating Multimodal Data for Joint Generative Modeling of Complex Dynamics

Abstract
Many, if not most, systems of interest in science are naturally described as nonlinear dynamical systems. Empirically, we commonly access these systems through time series measurements. Often such time series may consist of discrete random variables rather than continuous measurements, or may be composed of measurements from multiple data modalities observed simultaneously. For instance, in neuroscience we may have behavioral labels in addition to spike counts and continuous physiological recordings. While by now there is a burgeoning literature on deep learning for dynamical systems reconstruction (DSR), multimodal data integration has hardly been considered in this context. Here we provide such an efficient and flexible algorithmic framework that rests on a multimodal variational autoencoder for generating a sparse teacher signal that guides training of a reconstruction model, exploiting recent advances in DSR training techniques. It enables to combine various sources of information for optimal reconstruction, even allows for reconstruction from symbolic data (class labels) alone, and connects different types of observations within a common latent dynamics space. In contrast to previous multimodal data integration techniques for scientific applications, our framework is fully generative, producing, after training, trajectories with the same geometrical and temporal structure as those of the ground truth system.
Machine Learning, ICML, dynamical systems, chaos, generative models, scientific machine learning, recurrent neural networks, variational autoencoders, exploding gradient problem, teacher forcing
1 Introduction
For many temporally evolving complex systems in physics, biology, or the social sciences, we have only limited knowledge about the generating dynamical mechanisms. Inferring these from data is a core interest in any scientific discipline. It is also practically highly relevant for predicting important changes in system dynamics, like tipping points in climate systems (Bury et al., 2021; Patel & Ott, 2023). In recent years, a variety of machine learning (ML) methods for recovering dynamical systems (DS) directly and automatically from time series observations have been proposed (Brunton et al., 2016; Vlachas et al., 2018; Lusch et al., 2018; Pathak et al., 2018; Koppe et al., 2019b; Schmidt et al., 2021; Fu et al., 2019; Vlachas et al., 2020; Gauthier et al., 2021; Jordana et al., 2021; Lejarza & Baldea, 2022; Brenner et al., 2022; Hess et al., 2023; Chen & Wu, 2023; Yang et al., 2023; Vlachas et al., 2022), mostly based on recurrent neural networks (RNNs) for approximating the flow of the underlying true DS (Vlachas et al., 2020; Koppe et al., 2019b; Schmidt et al., 2021; Brenner et al., 2022; Hess et al., 2023; Vlachas et al., 2022). However, almost all of these methods assume that observed time series come as continuous signals with Gaussian noise from a single type of source. Yet, time series data from discrete random processes, like class labels or counts, are in fact quite commonplace in many areas, e.g., in the medical domain (electronic health records, smartphone-based data) (Koppe et al., 2019a), neuroscience (behavioral responses) (Schneider et al., 2023), or climate science (event counts) (Tziperman et al., 1997). Moreover, with the increasing availability of massive data acquisition techniques in many scientific disciplines, often one has simultaneous measurements from multiple different types of data channels with different statistical properties. For instance, in neuroscience one may have simultaneous recordings of spike events and calcium imaging together with behavioral class labels.
Multimodal data integration
In general, multimodal data integration is a thriving topic in many areas of artificial intelligence and ML research (Ahuja et al., 2017; Bhagwat et al., 2018; Baltrusaitis et al., 2019; Sutter et al., 2020; Shi et al., 2020; Antelmi et al., 2018; Liang et al., 2022; Lipkova et al., 2022; Xu et al., 2023; Steyaert et al., 2023; Warner et al., 2023; OpenAI, 2023). Generative models fusing multiple data channels produce a common latent code which allows for cross-modal prediction or output generation (e.g., images from language Amrani et al. (2021); Radford et al. (2021)), improves inference by complementing information too noisy or missing in one channel through recordings from other modalities (Qian et al., 2022), or may reveal interesting links among observed modalities (Liang et al., 2015).
Variational autoencoders (VAE) (Kingma & Welling, 2014; Rezende et al., 2014) are one popular variant of generative models which naturally lend themselves to multimodal settings (Baltrusaitis et al., 2019; Wu & Goodman, 2018; Sutter et al., 2020) and a sequential formulation (Bayer et al., 2021; Girin et al., 2021; Bai et al., 2021). Longitudinal autoencoders have been proposed (Ramchandran et al., 2021) to model temporal correlations in latent space, and have also been extended to multimodal data (Öğretir et al., 2022) for the purpose of time series forecasting (Antelmi et al., 2018; Bhagwat et al., 2018; Dezfouli et al., 2018; Shi et al., 2020; Sutter et al., 2020; Qian et al., 2022).
Dynamical Systems Reconstruction (DSR)
DSR, however, goes beyond mere forecasting in that we request a dynamical model of the data-generating process which captures the temporal and geometrical structure of the true underlying system (Koppe et al., 2019b; Abarbanel, 2022; Platt et al., 2023). A successful DS reconstruction can reveal important invariant topological and geometrical characteristics of the state space the DS lives in, its long-term temporal behavior (attractors), and its sensitivity to perturbations and parameter variations. It provides a mechanistic surrogate for the observed system which can be further analyzed, simulated, perturbed, or lesioned to obtain additional insight into its inner workings (Durstewitz et al., 2023). Often special training techniques (Mikhaeil et al., 2022; Brenner et al., 2022; Hess et al., 2023; Abarbanel, 2013, 2022; Vlachas & Koumoutsakos, 2023), regularization approaches (Schmidt et al., 2021; Platt et al., 2022; Jiang et al., 2023; Platt et al., 2023), or structural priors as in physics-informed neural networks (PINNs Raissi et al. (2019)), are required to achieve these goals. Plain ‘out-of-the-box’ gradient descent techniques or standard variational inference will often fail to reconstruct the underlying system’s long-term properties (Arribas et al., 2020; Brenner et al., 2022; Hess et al., 2023; Platt et al., 2023; Jiang et al., 2023; Durstewitz et al., 2023), as these standard procedures do not sufficiently constrain the vector field to prevent true and reconstructed trajectories from quickly diverging. This is especially a problem in chaotic systems, where it is related to the exploding or vanishing gradient problem in RNN training (Hochreiter & Schmidhuber, 1997; Hochreiter et al., 2001; Bengio et al., 1994; Mikhaeil et al., 2022; Hess et al., 2023). Control-theoretic techniques are often used to keep model-generated trajectories on track, or to synchronize them with those produced by the observed system, whilst training (Voss et al., 2004; Abarbanel et al., 2009; Abarbanel, 2013; Verzelli et al., 2021; Platt et al., 2021; Abarbanel, 2022). Similar ideas underlie the recently proposed methods of sparse (Mikhaeil et al., 2022; Brenner et al., 2022) and generalized (Hess et al., 2023) teacher forcing (STF and GTF, respectively) for learning chaotic DS, which balance model-generated and data-inferred states during training in a way that optimally controls gradient flows and trajectory divergence (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023).
Specific contributions
Although important in many areas of science, an efficient framework for multimodal data integration for the purpose of DS reconstruction is currently essentially lacking. A particular challenge here is to reconcile highly efficient training techniques for DS reconstruction like STF or GTF, which usually rest on an invertible decoder model, with the need to infer a common latent code from many different types of data. Here we address this challenge through a novel formulation of the multimodal data integration problem for DS reconstruction: Building on the success of variational autoencoders (VAEs) for multimodal integration (Baltrusaitis et al., 2019; Wu & Goodman, 2018; Sutter et al., 2020), we use multimodal VAEs (MVAEs) to construct a common latent representation from random variables following different distributional models. Crucially, however, rather than constructing a sequential VAE process that directly operates on this latent code (Kramer et al., 2022), here we employ the MVAE solely to create a multimodal teacher signal for guiding a DS reconstruction model in training. Special loss terms and shared decoder models ensure consistency between the MVAE’s and reconstruction model’s latent codes. This not only efficiently exploits the information available in multiple different data streams for DS reconstruction (as compared to a variety of alternative techniques), but for the first time, to our knowledge, enables DS reconstructions of chaotic systems from categorical information alone.
2 Multimodal Teacher Forcing (MTF) framework
Techniques like STF (Brenner et al., 2022; Mikhaeil et al., 2022) or GTF (Hess et al., 2023) enabled reconstructions even from challenging real-world data on which many previous methods failed. Hence, the key idea for making DS reconstruction from multiple, statistically distinct data sources work is the generation of a multimodal TF signal for efficient training of a reconstruction model. For this we utilize an MVAE which is then trained jointly with the DS reconstruction (DSR) model through a set of shared decoder models. This general framework, illustrated in Fig. 1, is very flexible and modular. In principle it can be used with any type of DSR model for approximating the flow of the observed system, as well as with any set of decoder and encoder models.
Figure 1: MTF setup. Multimodal observations are translated via an encoder into a common latent representation, which is used for sparse TF in the DSR model’s latent space. The latent trajectory is then mapped back into observation space via modality-specific decoder models.
DSR model
While completely flexible, for our specific experiments we chose the recently introduced ‘dendritic piecewise linear RNN (dendPLRNN)’ (Brenner et al., 2022) as a reconstruction model, as it is both expressive and has a mathematically tractable, piecewise linear design (Monfared & Durstewitz, 2020; Eisenmann et al., 2023) (other types of RNNs, such as LSTMs or GRUs, may be used and give similar results, see Suppl. Table B3). The dendPLRNN is defined by the $M$-dimensional latent process equation
$\bm{z}_{t}=$ $\bm{A}\bm{z}_{t-1}+\bm{W}\sum_{b=1}^{B}\alpha_{b}\max(\bm{0},\bm{z}_{t-1}-\bm{h}_{b})$ (1)
$+\bm{h}+\bm{U}\bm{s}_{t}+\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{\Sigma}),$
which describes the temporal evolution of an $M$-dimensional latent state vector $\bm{z}_{t}=(z_{1t}\dots z_{Mt})^{T}$ with diagonal matrix $\bm{A}\in\mathbb{R}^{M\times M}$, off-diagonal matrix $\bm{W}\in\mathbb{R}^{M\times M}$, noise covariance matrix $\bm{\Sigma}\in\mathbb{R}^{M\times M}$, a nonlinearity given by the ‘dendritic’ spline expansion with slopes $\alpha_{b}\in\mathbb{R}$ and thresholds $\bm{h}_{b}\in\mathbb{R}^{M}$ (Brenner et al., 2022), and a linear term $\bm{U}\in\mathbb{R}^{M\times P}$ weighing $P$-dimensional external inputs $\bm{s}_{t}$.
Decoder models
To infer this latent process equation jointly from multiple data modalities, the dendPLRNN is connected to different decoder models that take the distinct distributional properties of each modality into account. While our formulation is general and can work with any combination of continuous and/or discrete data types, for our exposition we consider as example time series of multivariate Gaussian ($\bm{X}$), ordinal ($\bm{O}$), [^1] Ordinal data, as often encountered in psychology or economics (Likert, 1932), are non-metric but still ordered. and count ($\bm{C}$) nature of length $T$, $\bm{Y}=\{\{\bm{x}_{1},\dots,\bm{x}_{T}\};\{\bm{o}_{1},\dots,\bm{o}_{T}\};\{\bm{c}_{1},\dots,\bm{c}_{T}\}\}$. In this case we may take as decoders a linear-Gaussian model, a cumulative link model for ordinal data, and a log-link function for Poisson data (see Methods A.1):
$\bm{x}_{t}|\bm{z}_{t}$ $\sim\mathcal{N}\left(\bm{B}\bm{z}_{t},\bm{\Gamma}\right);$ (2)
$\bm{o}_{t}|\bm{z}_{t}$ $\sim\operatorname{Ordinal}(\bm{\beta}\bm{z}_{t},\epsilon);$ (3)
$\bm{c}_{t}|\bm{z}_{t}$ $\sim\operatorname{Poisson}(\lambda(\bm{z}_{t})).$ (4)
Training: Multimodal Teacher Forcing (MTF)
Training RNNs on time series is generally challenging due to the exploding and vanishing gradient problem (Bengio et al., 1994). This becomes particularly severe in DS reconstruction as for chaotic systems exploding gradients cannot be avoided even in principle, due to the positive maximum Lyapunov exponent resulting in exponential divergence of trajectories (Mikhaeil et al., 2022; Hess et al., 2023). Yet, at the same time batch length cannot be too short, as the training algorithm will then fail to capture the system’s long term behavior and limit sets (Mikhaeil et al., 2022; Platt et al., 2022, 2023). STF (Mikhaeil et al., 2022) (or GTF (Hess et al., 2023)) therefore balance loss and trajectory divergence with the need to capture relevant long time scales in an optimal way. They do so by replacing forwarded-iterated latent states $\bm{z}_{t}$ by data-inferred states $\bm{\hat{z}}_{t}$ at strategically chosen time points (determined from the empirically estimated maximum Lyapunov exponent in the case of STF (Mikhaeil et al., 2022), or by averaging forward-propagated and data-inferred states in an optimal way in the case of GTF (Hess et al., 2023)). Data-inferred states are obtained by inverting the decoder model (e.g., $\bm{\hat{z}}_{t}=(\bm{B}^{T}\bm{B})^{-1}\bm{B}^{T}\bm{x}_{t}$ for a simple linear-Gaussian model as in eq. 2, i.e. taking the pseudo-inverse). This forcing then is only applied during model training, not at run time (see Methods A.1 for further details).
Inverting the decoder model is in general not possible, however, in particular for discrete random variables. Moreover, in the case of multiple simultaneously observed data modalities, it is unclear how to combine the different data modalities to obtain an optimal estimate $\bm{\hat{z}}_{t}$. We therefore need another means to create a (sparse) TF signal if we would like to utilize STF or GTF for training. Here we achieve this through an MVAE for building a joint latent representation over the different data types, $\bm{\tilde{z}}_{t}\sim p(\bm{\tilde{z}}_{t}|\bm{X},\bm{O},\bm{C})$, which can be used as a TF signal at time $t$. We denote the MVAE states by $\bm{\tilde{z}}_{t}\in\mathbb{R}^{K}$ to avoid confusion with the latent dynamical process $\bm{z}_{t}\in\mathbb{R}^{M}$ generated by the reconstruction model. [^2] Note that in general we do not assume the MVAE and reconstruction model latent codes to have the same dimensionality, i.e. $K\leq M$, to accommodate different forms of TF, see Methods A.1 for further details To ensure consistency between the DSR model’s and the MVAE’s latent states, the MVAE is coupled to the observations through the very same set of decoder models
$\bm{x}_{t}|\bm{\tilde{z}}_{t}$ $\sim\mathcal{N}\left(\bm{B}\bm{\tilde{z}}_{t},\bm{\Gamma}\right);$ (5)
$\bm{o}_{t}|\bm{\tilde{z}}_{t}$ $\sim\operatorname{Ordinal}(\bm{\beta}\bm{\tilde{z}}_{t},\epsilon);$ (6)
$\bm{c}_{t}|\bm{\tilde{z}}_{t}$ $\sim\operatorname{Poisson}(\lambda(\bm{\tilde{z}}_{t})),$ (7)
sharing all decoder parameters with the DSR model (see Fig. 1). The MVAE is trained by minimizing the negative Evidence Lower Bound (ELBO)
[…]
$+\log p_{\bm{\theta}}(\bm{\tilde{Z}})]-\mathbb{H}_{q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$
using the reparameterization trick for latent random variables (Kingma & Welling, 2014), where $\mathbb{H}_{q_{\bm{\phi}}}$ is the entropy of the approximate posterior, and assuming conditional independence of the observations given the latent states. We tested various choices for the encoder model ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$, achieving best results with a CNN (see Table B2 for results and Methods A.1 for architectural details).
Process prior and total loss
We thus have specified ${p_{\bm{\phi}}}(\bm{Y}|\bm{\tilde{Z}})$ through our choice of decoder models, and parameterized ${q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})$ by a CNN. But what is the best choice for the prior $p_{\bm{\theta}}(\bm{\tilde{Z}})$? Crucially, to further enforce consistency between the MVAE and DSR model latent codes, $\bm{\tilde{Z}}$ and $\bm{Z}$, respectively, we assume that $p_{\bm{\theta}}(\bm{\tilde{Z}})$ is given through the DSR model (the dendPLRNN in our case). Specifically, in line with the Gaussian assumptions in the dendPLRNN, eq. 1, we assume for the second term in eq. 8
$\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{\tilde{Z}})]\approx$ $\frac{1}{L}\sum_{l=1}^{L}\sum_{t=1}^{T}-\frac{1}{2}\bigl{(}\log|\bm{\Sigma}|$ (9)
$+(\bm{\tilde{z}}^{(l)}_{t}-\bm{\mu}_{t})^{\top}\bm{\Sigma}^{-1}(\bm{\tilde{z}}^{(l)}_{t}-\bm{\mu}_{t})$
$+const.\bigl{)},$
where the expectation value is approximated by $L$ Monte Carlo samples $\bm{\tilde{z}}^{(l)}_{t}\sim q_{\bm{\phi}}(\bm{\tilde{z}}_{t}|\bm{Y})$, and the means $\bm{\mu}_{t}=\mathbb{E}(\bm{z}_{t}|\bm{z}_{t-1})$ directly come from the DSR model (see Methods A.1 for details). We therefore also call this term, eq. 9, the consistency loss, $\mathcal{L}_{\text{con}}$.
The total training loss is now given by the MVAE’s negative ELBO, eq. 8, combined with the reconstruction loss (negative log-likelihood) of the DSR model (funneled through the same set of decoder models), i.e. by adding the term $-\log p_{\bm{\theta}}(\bm{Y}|\bm{Z})=:\mathcal{L}_{\text{DSR}}$ to eq. 8, where $\bm{Z}$ are the latent states of the DSR model:
$\mathcal{L}_{\text{MTF}}=-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})]-\mathbb{H}_{q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})+\mathcal{L}_{\text{con}}+\mathcal{L}_{\text{DSR}}$ (10)
Whilst training, we then use $\bm{\tilde{z}}_{t}$ as our TF signal for guiding the DSR model. Specifically, for STF we replace (a subset of the) latent states $\bm{z}_{t}$ by $\bm{\tilde{z}}_{t}$ at strategically chosen times $l\tau+1,\ l\in\mathbb{N}_{0}$, as illustrated in Fig. 1 (similarly, for GTF we may take a weighted average; see (Hess et al., 2023) for details). Note that our usage of an MVAE here thus profoundly differs from the more conventional use in, for instance, sequential VAEs (Sutter et al., 2020; Kramer et al., 2022). Ablation studies further confirm that all components of the loss (Fig. A3), as well as an optimal choice for the TF interval (Fig. A5), are crucial to the success of the approach. Specifically, optimal TF intervals were associated with smooth and well navigable loss landscapes (Fig. A4).
3 Results
3.1 Evaluating MTF: Comparison to other algorithmic strategies
We first compared MTF on moderately challenging multimodal DS reconstruction tasks [^3] Moderately challenging because many of the comparison methods actually broke down on the more difficult setups. to a variety of other possible algorithmic strategies that may be considered, all of which lacking one or more key ingredients of our methodology. For this, a training and a test set of $100,000$ time steps were created from a Lorenz-63 and a Rössler system with $1\%$ process noise, and a $6$d Lewis-Glass network model (Lewis & Glass, 1992; Gilpin, 2022), all in their chaotic regimes (see Methods A.5 for detailed parameter settings and numerical integration). From the simulated trajectories we then sampled ordinal and count observations using Eqs. (13) and (16) (with randomly drawn parameters), as well as continuous observations with $10\%$ Gaussian noise. Example reconstructions by MTF are in Fig. 2(a) and Fig. A9.
We compared the performance of MTF on these simulated data to the following setups: First, to a sequential multimodal VAE ((Kramer et al., 2022); to our knowledge the only other general approach intended for DSR from multimodal data). Second, we tried classical RNN training, where observations are provided as inputs at every time step, via truncated BPTT using modality specific decoder models, Eqs. 2-4. A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’, a method suggested in the dynamical systems literature (Voss et al., 2004), but to our knowledge not extended to multimodal data so far. We also included two other comparisons where we first transformed multimodal data to be approximately Gaussian (via Box-Cox & Gaussian kernel smoothing, see Methods A.3), and then either trained the RNN via standard BPTT-TF (Brenner et al., 2022) or by VAE-TF, similar as in MTF, but without modality-specific decoder models (labeled GVAE-TF). For comparability the same RNN architecture (the dendPLRNN, eq. 1) was used in all these comparisons.
Since in DS reconstruction we are primarily interested in invariant and long-term properties of the underlying DS (Mikhaeil et al., 2022; Platt et al., 2023; Jiang et al., 2023; Hess et al., 2023; Platt et al., 2021), for evaluation we focused on measures that capture the geometrical structure in state space ($D_{\text{stsp}}$, a type of Kullback-Leibler divergence (Koppe et al., 2019b; Brenner et al., 2022; Hess et al., 2023)) and the asymptotic temporal structure (evaluated either through the Hellinger distance, $D_{\text{H}}$, on power spectra (Mikhaeil et al., 2022; Hess et al., 2023), or auto-covariance functions for the ordinal and categorical data; Methods A.2). We also report mean-squared errors (MSE) to assess short-term ahead prediction ($10$-step-ahead $L_{2}$ prediction error (PE) for continuous and $L_{1}$ PE for ordinal data (Öğretir et al., 2022)). However, especially in chaotic systems, PEs are not suited for evaluating the system’s longer-term behavior because of the exponential divergence of nearby trajectories (Wood, 2010; Mikhaeil et al., 2022). Details of all measures are provided in Methods A.2. As evidenced in Table 1, MTF outperforms all other possible model setups, sometimes by large margins.
Table 1: Comparison of dendPLRNN trained by MTF (proposed method), by a sequential multimodal VAE (SVAE) based on (Kramer et al., 2022), a VAE-TF approach similar to MTF except that all data modalities were ‘Gaussianized’ (GVAE-TF), BPTT-TF as in (Brenner et al., 2022) using Gaussianized data, and a multiple-shooting (MS) approach. Training was performed on multivariate normal, ordinal, and count data produced by the chaotic Lorenz system, Rössler system, and Lewis-Glass model. Values are mean $\pm$ SEM, averaged over 15 trained models. X = value cannot be computed for this model (e.g., because resp. decoder model is not present). Note that SCC (Spearman cross-correlation), OACF (ordinal autocorrelation function), and CACF (count autocorrelation function) all refer to mean-squared-errors (MSEs) between ground truth and generated correlation functions. Bold numbers indicate top performance within ±$\pm 1$ SEM.
Dataset Method $D_{stsp}$downarrow $D_{H}$downarrow PE $\downarrow$ OPE $\downarrow$ SCC $\downarrow$ OACF $\downarrow$ CACF $\downarrow$
Lorenz MTF $\mathbf{3.4\pm 0.35}$ $\mathbf{0.30\pm 0.06}$ $\mathbf{1.3\mbox{e$-$}2}\pm\mathbf{2\mbox{e$-$}4}$ $\mathbf{0.12\pm 0.03}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{6.6\mbox{e$-$}5\pm 8.1\mbox{e$-$}6}$
SVAE $11.1\pm 0.6$ $0.82\pm 0.05$ ${6.3\mbox{e$-$}1}$$\pm$ ${5.1\mbox{e$-$}2}$ $0.68\pm 0.03$ $0.14\pm 0.01$ $0.18\pm 0.02$ ${8.5\mbox{e$-$}5}$$\pm$ ${1.6\mbox{e$-$}5}$
BPTT $6.31\pm 1.2$ $0.47\pm 0.11$ ${2.1\mbox{e$-$}1}$$\pm$ ${2.4\mbox{e$-$}2}$ $0.33\pm 0.04$ $0.11\pm 0.02$ $0.09\pm 0.02$ ${8.2\mbox{e$-$}5}$$\pm$ ${9\mbox{e$-$}6}$
MS $4.5\pm 1.5$ $0.61\pm 0.08$ X X $0.14\pm 0.04$ $0.11\pm 0.02$ $\mathbf{{6.5\mbox{e$-$}5}\pm{3.8\mbox{e$-$}6}}$
GVAE-TF $4.3\pm 0.3$ $0.47\pm 0.07$ ${3.6\mbox{e$-$}1}\pm{1.5\mbox{e$-$}3}$ X X X X
BPTT-TF $8.8\pm 1.9$ $0.86\pm 0.05$ ${4.4\mbox{e$-$}1}\pm{2.2\mbox{e$-$}2}$ X X X X
Rössler MTF $\mathbf{1.45\pm 0.71}$ $\mathbf{0.32\pm 0.03}$ $\mathbf{{1.9\mbox{e$-$}3}\pm{7.1\mbox{e$-$}5}}$ $\mathbf{0.08\pm 0.02}$ $\mathbf{0.04\pm 0.004}$ $\mathbf{0.017\pm 0.003}$ $\mathbf{{6.5\mbox{e$-$}5}\pm{1.2\mbox{e$-$}5}}$
SVAE $10.7\pm 1.5$ $0.66\pm 0.05$ ${1.5\mbox{e$-$}1}$$\pm$ ${3.1\mbox{e$-$}2}$ $0.24\pm 0.02$ $0.17\pm 0.03$ $0.13\pm 0.02$ ${1.1\mbox{e$-$}4}\pm{1.4\mbox{e$-$}5}$
BPTT $9.05\pm 0.5$ $0.7\pm 0.01$ ${7.4\mbox{e$-$}2}$$\pm$ ${2.0\mbox{e$-$}3}$ $0.18\pm 0.02$ $0.3\pm 0.03$ $0.19\pm 0.07$ ${1.5\mbox{e$-$}4}$$\pm$ ${6\mbox{e$-$}6}$
MS $3.99\pm 1.1$ $0.59\pm 0.04$ X X $0.08\pm 0.04$ $0.09\pm 0.02$ ${1.6\mbox{e$-$}4}$$\pm$ ${5.9\mbox{e$-$}5}$
GVAE-TF $12.1\pm 0.5$ $0.55\pm 0.04$ ${4.9\mbox{e$-$}2}\pm{3.4\mbox{e$-$}3}$ X X X X
BPTT-TF $8.9\pm 1.4$ $0.64\pm 0.07$ ${2.8\mbox{e$-$}1}\pm{1.8\mbox{e$-$}3}$ X X X X
Lewis-Glass MTF $\mathbf{0.27\pm 0.07}$ $\mathbf{0.33\pm 0.02}$ $\mathbf{{2.1\mbox{e$-$}3}\pm{7\mbox{e$-$}5}}$ $\mathbf{0.11\pm 0.01}$ ${0.12\pm 0.03}$ $\mathbf{0.05\pm 0.02}$ ${2.3\mbox{e$-$}4}\pm{2.0\mbox{e$-$}5}$
SVAE $2.6\pm 0.5$ $0.52\pm 0.03$ ${8.0\mbox{e$-$}2}$$\pm$ ${4\mbox{e$-$}3}$ $0.26\pm 0.01$ $0.4\pm 0.05$ $0.18\pm 0.03$ ${7.5\mbox{e$-$}3}\pm{4.7\mbox{e$-$}3}$
BPTT $2.8\pm 0.5$ $0.57\pm 0.05$ ${6.2\mbox{e$-$}2}$$\pm$ ${3\mbox{e$-$}3}$ $0.23\pm 0.02$ $0.47\pm 0.08$ $0.21\pm 0.03$ ${9.1\mbox{e$-$}3}\pm{3.2\mbox{e$-$}3}$
MS $\mathbf{0.33\pm 0.06}$ $\mathbf{0.35\pm 0.01}$ X X $\mathbf{0.08\pm 0.01}$ $\mathbf{0.04\pm 0.002}$ $\mathbf{{1.9\mbox{e$-$}4}\pm{7.5\mbox{e$-$}6}}$
GVAE-TF $\mathbf{0.28\pm 0.08}$ $0.44\pm 0.02$ ${4.6\mbox{e$-$}3}\pm{4\mbox{e$-$}4}$ X X X X
BPTT-TF $2.51\pm 0.71$ $0.43\pm 0.03$ ${2.6\mbox{e$-$}2}\pm{3\mbox{e$-$}3}$ X X X X
To investigate multimodal data integration in a more challenging setting, we next tested a scenario where continuous observations from the Lorenz-63 system were heavily distorted by Gaussian noise with $50\%$ of the data variance. At the same time, ordinal observations with $8$ variables divided into $7$ ordered categories each, $o_{nt}\in\{1\dots 7\}$, $n=1\dots 8$, were sampled using eq. 13. We then trained the dendPLRNN via MTF once with, and once without, ordinal observations. Fig. 2(b) proves that with ordinal observations on board, DS reconstruction is, in principle, possible even under these challenging conditions. The cumulative histograms of the geometric measure $D_{stsp}$ and the temporal measure $D_{H}$ in Fig. 2(c), furthermore show that inclusion of ordinal observations significantly improves reconstructions.
Figure 2: DS reconstruction from moderately (a) and heavily (b-c) distorted continuous observations (Gaussian observation noise of $10\%$ and $50\%$, respectively, of the data variance) and other simultaneously provided observation modalities, sampled from a Lorenz-63 system. (a) Freely generated example trajectories from a dendPLRNN ($M=20,B=10,K=20,\tau=10$) trained with MTF jointly on Gaussian ($10\%$ noise), ordinal, and count data. (b) Same as (a) for a dendPLRNN trained by MTF on heavily distorted Gaussian ($50\%$ noise) and ordinal observations. Note that even in this case the butterfly wing structure of the Lorenz attractor was successfully captured. The maximum Lyapunov exponent ($\lambda_{\text{max}}$) furthermore confirms the dendPLRNN-generated attractor is chaotic (for the GT Lorenz system, $\lambda_{\text{max}}\approx 0.903$). (c) Normalized cumulative histograms of geometrical attractor disagreement ($D_{stsp}$, left) and Hellinger distance ($D_{H}$, right) between reconstructed and ground-truth system for the same setting as in (b).
3.2 Dynamical systems reconstruction from discrete random variables
Motivated by these results, we pushed the system even further and attempted DS reconstruction solely from ordinal data (created as above), i.e. completely omitting continuous observations. This is profoundly more challenging than the multimodal setting with Gaussian data, since the ordinal process considerably coarse-grains the underlying continuous dynamical process. Since in this case we do not have a direct linear mapping between ground truth state space and that of the trained dendPLRNN (which in the case of Gaussian observations would simply be given by the linear decoder, eq. 11), we construct one post-hoc by optimizing a linear operator given by a linear dimensionality reduction (PCA) concatenated with a geometry-preserving rotation operation (see Methods A.2). Fig. 3 (top) confirms that, using MTF, successful DS reconstruction is still possible under these conditions.
Finally, we asked whether DS reconstruction might be feasible based on a purely symbolic representation of the dynamics. For this, $4^{3}$ symbols were used corresponding to subregions defined by a $4\times 4\times 4$ grid superimposed on the attractor (this symbolic code was reduced further and then delay-embedded, see Suppl. A.5 for full details). Fig. 3 (bottom) illustrates the general procedure, and also shows that, using MTF, successful DS reconstruction is – in principle – possible from just a symbolic (categorical) coding of the underlying chaotic attractor. This is to our knowledge the first time such a result has been shown. Comparable results could not be achieved by the multimodal sequential VAE or MS (see Table B4). Fig. A8 shows that even the maximum Lyapunov exponent $\lambda_{max}$ of the system reconstructed with MTF from just the symbolic coding of the Lorenz attractor often comes close to the true value (see also Fig. A6).
Figure 3: DS reconstruction from discrete observations by MTF ($M=30,B=15,K=30,\tau=10$). Top: Reconstruction of Rössler attractor from only ordinal time series. Bottom: DS reconstruction from symbolic coding of Lorenz attractor (see Fig. A7 for true and predicted class label probabilities). Note that in both cases the topology and general geometry are preserved, and maximal Lyapunov exponents closely match those of the true systems (Rössler: $\lambda_{\text{max}}^{\text{true}}\approx 0.072$, Lorenz: $\lambda_{\text{max}}^{\text{true}}\approx 0.903$). TDE = temporal delay embedding.
3.3 Reconstruction from neuroimaging and behavioral data
We evaluated MTF’s performance on real multimodal time series using a dataset of functional magnetic resonance imaging (fMRI) recordings from human subjects undertaking various cognitive tasks in an fMRI scanner (Koppe et al., 2014). The data consisted of continuous blood-oxygenation-level-dependent (BOLD) time series acquired from 26 subjects at a sampling rate of $1/3$ $\mathrm{Hz}$. The first principal component of BOLD activity from each of 20 different brain regions was selected for analysis. The individual time series were rather short ($T=360$ per subject), posing a significant challenge for DS reconstruction. In addition to continuous BOLD activity, categorical time series of cognitive stage labels were given for each subject, corresponding to the five cognitive tasks that each subject went through repeatedly during the experiment (‘Rest’, ‘Instruction’, ‘Choice Reaction Task [CRT]’, ‘Continuous Delayed Response Task [CDRT]’ and ‘Continuous Matching Task [CMT]’) [^4] Note that CRT, CDRT, and CMT did not differ in terms of presented stimuli or response types, but only in cognitive load, thus reflecting processes internal to the subject., which we accommodated via a multi-categorical decoder model, eq. 15 in Methods A.1.
[…]
Table 2: Comparison among multi-modal reconstruction methods for experimental fMRI+behavioral data. For each subject and training method, medians across $15$ trained models were first obtained for each measure, which were then averaged across $20$ subjects ($\pm$ SEM). SEM = standard error of the mean. X = value not accessible for this method. For abbreviations see Table 1.
Dataset Method $D_{stsp}$downarrow $D_{H}$downarrow PE $\downarrow$
fMRI MTF $\mathbf{0.55\pm 0.04}$ $\mathbf{0.301\pm 0.007}$ $\mathbf{1.21\pm 0.08}$
SVAE $1.9\pm 0.22$ $0.441\pm 0.019$ $2.34\pm 0.12$
BPTT $3.31\pm 0.8$ $0.52\pm 0.05$ $2.8\pm 0.15$
MS $1.06\pm 0.14$ $0.373\pm 0.012$ X
GVAE-TF $0.67\pm 0.06$ $0.335\pm 0.011$ $1.64\pm 0.07$
BPTT-TF $0.63\pm 0.03$ $0.312\pm 0.006$ $1.39\pm 0.05$
Fig. A2(b) further shows an example where on a left-out test set simulated BOLD activity and model-generated task labels closely resembled those of the real subject. Fig. A2(c) investigates to what extent cross-modal links have been built by the trained system in its latent space by illustrating that after multimodal training, freely sampled latent activity trajectories from the dendPLRNN reflect the structure of the different tasks stages. Fig. A10 in Suppl. A.5 provides generated time series for several example subjects. Table 2, finally, compares the DSR performance on the fMRI and behavioral data for all the methods introduced in sect. 3, with MTF substantially outperforming the sequential MVAE from (Kramer et al., 2022).
3.4 DSR from neurophysiological and location data
As a second empirical test case we used electrophysiological recordings of multiple single neurons from the CA1 region of the rodent hippocampus (Grosmark et al., 2016), a brain region involved in spatial navigation and memory (Buzsaki, 2006; Moser et al., 2017). During the experiments the rats traversed a 1.6 m long track along which the longitudinal position of the animal was recorded, treated as a continuously valued Gaussian $1$d time series. Water rewards were provided at each of the two end points of the track, indicated to the model through brief external inputs. At the same time, for the data set on which we focused our analysis, the spiking activity (action potential time points) of $120$ neurons was recorded via intracortical electrodes (of which $60$ with reasonably high spiking rate were retained). Following common procedures in neuroscience (Zhou & Wei, 2020; Schneider et al., 2023), spike trains were segmented into $200$ ms bins, resulting in a multivariate count time series (for details, see Methods A.5). These were modelled with a negative-binomial decoder (capturing the high dispersion better than Poisson or zero-inflated Poisson models).
Multimodal (position and spike count) data were split into a training and a test set of $4600$ time steps each. Example true and model-generated spike trains for the test set are in Fig. 4 (a) (see Fig. A11 for training set). Fig. 4 (b) compares various standard spike train statistics like the mean firing rate, mean zero rate, coefficient of variation, and the neural cross-correlation matrix between experimental and model-generated data, iterated forward from one inferred initial state across the whole test set. The agreement in spiking statistics between model-simulated long-term behavior and test data was in fact as good as between different segments (training and test set) of the experimental data itself. Moreover, predicted position on the track closely matched that of the real behavior for the test set (Fig. 4 (a)). Using both data sources (spike trains and position) led to far better reconstructions across all spike train performance metrics, and overall more robust reconstructions, than when only spike count data were provided for model training (Mann-Whitney U-test across $50$ trained models, $p<0.025$ for all metrics; Fig. 4 (c)). In the CA1 region, place cells are known to encode spatial information (O’Keefe & Nadel, 1978). Our results indicate that the MTF algorithm can leverage this information to build a joint latent model of the neural recordings and movement of the rat (Fig. 4 (d)).
Figure 4: (a) Example reconstructions of spike trains and spatial location of animal. Top: Spike train data from a test set not used for model training (topmost), and model-generated spike trains (below) simulated from a data-inferred initial condition. Bottom: True and model-predicted spatial position on the test set. (b) Top: Correlation of various spike train statistics between test set and model-generated data (blue), and – for comparison – between experimental training and test set data (orange): mean spike rate, zero count ratio, coefficient of variation, and correlation between cross-correlation coefficients. Diagonal gray lines are bisectrices, not regression lines. Bottom: Cross-correlation matrices among all neurons for the experimental test set (left) and model-generated spike data (right). (c) Joint DSR from both spatial and neural data significantly improves reconstructions compared to just neural data alone. (d) DSR model latent space (shown is a subspace), illustrating how the latent dynamics is organized according to the animal’s spatial position (color-coded).
4 Conclusions
Here we introduced a novel training framework (MTF) for DSR from non-Gaussian and multimodal time series data. While DS reconstruction is meanwhile a large field in scientific machine learning (Brunton et al., 2016; Pathak et al., 2018; Lejarza & Baldea, 2022; Brenner et al., 2022; Hess et al., 2023; Chen & Wu, 2023), reconstruction based on discrete or multimodal data has hardly been addressed so far, although such scenarios are commonplace in many areas like medicine, neuroscience, or climate research. Here we utilize recent insights on guiding the training process by control signals, as in STF (Mikhaeil et al., 2022) and GTF (Hess et al., 2023), within a novel multimodal DSR framework. In our approach, a sparse TF signal is generated by an MVAE that translates different data modalities into a common latent code in order to guide a DSR model. Our general framework is very flexible and allows to easily swap DSR, encoder, or decoder models for the problem at hand. For diverse benchmarks and sampling conditions, MTF clearly outperforms various other training setups and models. We further highlighted on examples from neuroscience how the DSR model’s latent space exposes meaningful relations between different data modalities (neural activity and behavior in our case). In contrast to other recently proposed methods for construing a common latent space from multiple data sources (Schneider et al., 2023), our method actually delivers a generative model of the system dynamics. A crucial difference to ‘classical’ multimodal sequential VAEs (Kramer et al., 2022) is that our algorithm allows to optimally control trajectory and gradient flows whilst training, a key ingredient for successful DS reconstruction (Mikhaeil et al., 2022; Hess et al., 2023; Eisenmann et al., 2023).
This way, MTF was even able, under certain conditions, to recover chaotic attractors from ordinal or categorical observations alone, including sensible estimates of their maximum Lyapunov exponent. This, to our knowledge, has never been shown before, and at first glance appears puzzling, as an ordinal or symbolic coding of the dynamics seems to remove much of the geometrical information. However, results in the field of symbolic dynamics ensure that under certain conditions a symbolic encoding of a DS preserves topological invariants and Lyapunov spectra (Hubler, 2012; Osipenko, 2000; McCullough et al., 2017; Zhang et al., 2017), and delay embedding theorems have been formulated for non-continuous signals like point processes (Sauer, 1994, 1995). While here we presented first results on DS reconstruction for discrete data, whether and how much of the original state space topology of a data-generating DS can be recovered from non-continuous, non-Gaussian random variables remains an important topic for future theoretical and empirical research.
Acknowledgements
This work was funded by the European Union’s Horizon 2020 research and innovation Programme under grant agreement 945263 (IMMERSE), by the German Research Foundation (DFG) within the collaborative research center TRR-265 (project A06), by the German Research Foundation (DFG) within Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and by a living lab grant by the Federal Ministry of Science, Education and Culture (MWK) of the state of Baden-Württemberg, Germany (grant number 31-7547.223-7/3/2), to DD and GK.
Ethics Statement
This paper presents a novel approach in the context of scientific machine learning, with applications in fields like neuroscience. We believe the primary ethical implications of this approach are positive. While we can not specifically highlight any ethical concerns, given the broad scope of possible applications of our approach, potential misuses can not be a priori excluded.
Appendix A Methods
A.1 MTF framework
The MTF framework consists of four major components, 1) a DSR model for approximating the flow of the underlying, observed DS, 2) a set of modality-specific decoder models which capture the statistical particularities of the observed time series, 3) a MVAE for integrating all different data modalities in order to produce a joint latent code which can be used for TF, and 4) the actual MTF training algorithm. Here we describe each of these components in turn.
DSR model
While our framework is generic and allows to work with any type of DSR model, here we used the dendPLRNN, eq. 1, recently introduced specifically for DSR (Brenner et al., 2022) (see Suppl. Table B3 for comparison with other RNN models). A major advantage of the dendPLRNN and related formulations is that it provides some degree of analytical tractability, which is of major importance in scientific settings where we are usually interested in analyzing the data-inferred model further as a formal surrogate for the real system. For example, for the dendPLRNN we have efficient algorithms for precisely locating all fixed points, cycles, and bifurcation manifolds (Eisenmann et al., 2023), and it allows for translation into an equivalent continuous-time representation (Monfared & Durstewitz, 2020), which further eases certain types of DS analysis.
Decoder models
To infer the latent process dendPLRNN jointly from multiple data modalities, it is connected to different decoder models that take the distinct distributional properties of each modality into account, Eqn. 2-4.
For normally distributed data, the decoder may take the simple linear Gaussian form
$\bm{x}_{t}=\bm{B}\bm{z}_{t}+\bm{\eta}_{t},$ (11)
with factor loading matrix $\bm{B}\in\mathbb{R}^{N\times M}$, and Gaussian observation noise $\bm{\eta}_{t}\sim\mathcal{N}(\bm{0},\bm{\Gamma})$ with covariance $\bm{\Gamma}\in\mathbb{R}^{N\times N}$ (assumed to be diagonal here).
Ordinal data, in contrast, are not associated with a metric space, but there is a natural ordering between variables, as e.g. in survey data in economy or psychology commonly assessed through Likert scales (Likert, 1932). Treating ordinal data as metric can lead to a variety of problems, as pointed out in (Liddell & Kruschke, 2018). Ordinal observations are coupled to latent states via a generalized linear model (McCullagh, 1980). Here, specifically, we assume that the ordinal observations $\bm{o}_{t}$ are derived from an underlying unobserved continuous variable $u_{it}$, which is linked to the latent states $\bm{z}_{t}$ via a linear model
$u_{it}=\bm{\beta}_{i}^{T}\bm{z}_{t}+\epsilon_{it},$ (12)
where $\bm{\beta}_{i}^{T}\in\mathbb{R}^{M}$ are the model parameters and $\epsilon_{it}$ is an independently distributed noise term. The distributional assumptions about the noise term $\epsilon_{it}$ determine which link function to use. A Gaussian assumption leads to an ordered probit model, while a logistic assumption leads to an ordered logit model (Winship & Mare, 1984). While both models lead to similar results, we found the ordered logit model to work slightly better in practice, and hence we focus on it here. Inverting the link function leads to an expression for the cumulative probabilities:
$p\left(o_{it}\leq k|\bm{z}_{t}\right)=\frac{\exp\left(\beta_{ik}^{0}-\bm{\beta}_{i}^{T}\bm{z}_{t}\right)}{1+\exp\left(\beta_{ik}^{0}-\bm{\beta}_{i}^{T}\bm{z}_{t}\right)}.$ (13)
The probability masses $p\left(o_{it}=k|\bm{z}_{t}\right)$ follow from the cumulative distribution via $p\left(o_{it}=k|\bm{z}_{t}\right)=p\left(o_{it}\leq k|\bm{z}_{t}\right)-p\left(o_{it}\leq k-1|\bm{z}_{t}\right)$, from which we can compute the log-likelihood as
$\log p_{\bm{\theta}}(\bm{O}|\bm{Z})=\sum_{i}^{N}\sum_{t}^{T}\sum_{k}^{K}\left[o_{it}=k\right]\log p\left(o_{it}=k|\bm{z}_{t}\right).$ (14)
Categorical observations are, like ordinal observations, not associated with a metric space, and in contrast to ordinal data there is also no natural ordering between variable values. To couple categorical observations to the latent states, we employed the natural link function given by
$\pi_{i}=\frac{\exp\left(\bm{\beta}_{i}^{T}\mathbf{z}_{t}\right)}{1+\sum_{j=1}^{K-1}\exp\left(\bm{\beta}_{j}^{T}\mathbf{z}_{t}\right)}\quad\forall i\in\{1\ldots K-1\}$ (15)
$\pi_{K}=\frac{1}{1+\sum_{j=1}^{K-1}\exp\left(\bm{\beta}_{j}^{T}\mathbf{z}_{t}\right)}$
Here, the parameters $\bm{\beta}_{i}\in\mathbb{R}^{M\times 1}$ constitute the respective regression weights for category $i=1\ldots K-1$, with the total probability over all categories $\sum_{i=1}^{K}\pi_{i}=1$.
For count observations $\left\{\bm{c}_{t}\right\}_{t=1}^{T}$, with $\bm{c}_{t}=\left(c_{1t},\dots,c_{Lt}\right)^{T}$, we tried two different decoder models. First, a standard Poisson model,
$\quad p_{\bm{\theta}}\left(c_{lt}|\bm{z}_{t}\right)=\frac{\lambda_{lt}^{c_{lt}}}{c_{lt}!}e^{-\lambda_{lt}}.$ (16)
These probabilities are related to the latent states via the log-link function, $\log\lambda_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, where $\bm{\gamma}^{(l)}$ is a vector of coefficients. Thus, $\lambda_{lt}=e^{\gamma_{0}^{(l)}+\bm{\gamma}^{(l)}\bm{z_{t}}}$ is the expected count for the $l^{\text{th}}$ observation at time $t$.
Second, a negative binomial model given by
$p_{\bm{\theta}}\left(c_{lt}|\bm{z}_{t}\right)=\frac{\Gamma(c_{lt}+\phi_{l})}{\Gamma(\phi_{l})c_{lt}!}\left(\frac{\phi_{l}}{\mu_{lt}+\phi_{l}}\right)^{\phi_{l}}\left(\frac{\mu_{lt}}{\mu_{lt}+\phi_{l}}\right)^{c_{lt}},$ (17)
where $\mu_{lt}$ is the mean count and $\phi_{l}$ the dispersion parameter of the negative binomial distribution for the $l^{\text{th}}$ observation at time $t$. As for the Poisson model, we use a log-link function, $\log\mu_{lt}=\gamma_{0}^{(l)}+\sum_{m=1}^{M}\gamma_{m}^{(l)}z_{mt}$, with $\bm{\gamma}^{(l)}$ a vector of coefficients, and $z_{mt}$ the $m^{\text{th}}$ latent variable at time $t$. Properly accounting for dispersion significantly improved modeling of the spike counts. The likelihood function is based on the torch.distributions.NegativeBinomial class from the PyTorch library.
Multimodal variational autoencoder
The MVAE consists of a set of decoder models, $p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})$, shared with the DSR model, with examples given in Eqs. 5-7 and above, a prior model $p_{\bm{\theta}}(\bm{\tilde{Z}})$ instantiated by the DSR model, see eq. 9, and the encoder (approximate posterior) $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$, where a sample from this distribution provides the TF signal (see also Fig. 1). This setup is modular and all model components may be flexibly chosen. We assume conditional independence (given latent states) of the observations, such that the likelihood terms related to the observations simply sum up, i.e. $\log p_{\bm{\theta}}(\bm{{Y}}|\bm{\tilde{Z}})=\sum_{t=1}^{T}\left(\log p_{\bm{\theta}}(\bm{x}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{o}_{t}|\bm{\tilde{z}}_{t})+\log p_{\bm{\theta}}(\bm{{c}_{t}}|\bm{\tilde{z}}_{t})\right)$ for the decoders specified in Eqs. 5-7 (with parameters shared with Eqs. 2-4).
For the encoder part, as the latent dendPLRNN process itself is conditionally Gaussian, we make a Gaussian assumption for the variational density $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})=\mathcal{N}(\bm{\mu}_{\bm{\phi}}(\bm{Y}),\bm{\Sigma}_{\bm{\phi}}(\bm{Y}))$ to approximate the true posterior $p(\bm{\tilde{Z}}|\bm{Y})$, where mean and covariance are functions of the data. We further use the common mean field approximation to factorize $q_{\bm{\phi}}(\bm{\tilde{Z}}|\bm{Y})$ across time (Girin et al., 2021). The mean $\bm{\mu}_{\theta}(\bm{Y})$ and covariance $\bm{\Sigma}_{\theta}(\bm{Y})$ of the approximate posterior are parameterized by one-dimensional (across time) convolutional layers (Brenner et al., 2022), using observations within the window defined by the input kernel. For the mean we use a 4-layer stack of convolutional layers with decreasing kernel sizes $11,8,5,3$, while the diagonal covariance was parameterized by a single convolutional layer with a kernel size of $11$. We also tested various other choices for the encoder, as discussed in Suppl. A.4.
Multimodal teacher forcing
Once we have a mechanism for generating a multimodal TF signal, our framework can accommodate any form of TF training scheme, like STF (Mikhaeil et al., 2022) or GTF (Hess et al., 2023). Consider an observed (multimodal) time series $\bm{Y}=\{\bm{y}_{1},\bm{y}_{2},\dots,\bm{y}_{T}\}$ generated by a DS we would like to reconstruct. From this we create a control series $\hat{\bm{Z}}=\{\hat{\bm{z}}_{1},\hat{\bm{z}}_{2},\dots,\hat{\bm{z}}_{T}\}$, $\hat{\bm{z}}_{t}\in\mathbb{R}^{M}$. In STF or GTF, one obtains $\hat{\bm{Z}}$ from inverting the decoder model, projecting observations into the DSR model’s latent space. Here, in contrast, we equate the control series with the latent code $\bm{\tilde{Z}}=\{\bm{\tilde{z}}_{1},\dots,\bm{\tilde{z}}_{T}\}$, $\bm{\tilde{z}}_{t}\in\mathbb{R}^{K}$, produced by the MVAE. For the present evaluation we focused on a specific form of STF (Brenner et al., 2022), where the first $K$ latent states are replaced by the control states $\hat{z}_{k,l\tau+1}=\tilde{z}_{k,l\tau+1},\ k\leq K$, at times $l\tau+1,\ l\in\mathbb{N}_{0}$, with forcing interval $\tau\geq 1$, while the remaining latent states, $\hat{z}_{k,l\tau+1}=z_{k,l\tau+1},\ k>K$, remain unaffected. Defining =$\mathcal{F}=\{l\tau+1\}_{l\in\mathbb{N}_{0}}$, the dendPLRNN updates can thus be written as
$\bm{z}_{t+1}=\begin{cases}\text{dendPLRNN}(\bm{\hat{z}}_{t})&\text{if $t\in\mathcal{F}$}\\ \text{dendPLRNN}(\bm{z}_{t})&\text{else}\end{cases}$ (18)
For the initial condition $\bm{z}_{1}$ of the latent process, if $K<M$, the $M-K$ remaining states are randomly sampled from a standard normal distribution. As shown in Mikhaeil et al. (2022), the best tradeoff between exploding gradients and capturing relevant long-term dependencies is achieved when choosing the forcing interval $\tau$ according to the system’s maximal Lyapunov exponent (predictability time). However, for non-continuous data (like counts), this cannot readily be determined, in which case the forcing interval $\tau$ may be regarded as a hyper-parameter determined via grid search (see also Suppl. Fig. A5 for example reconstructions for different choices of $\tau$).
To ensure that the dendPLRNN and MVAE can share all decoder model parameters, only the first $K\leq M$ states of the generated latent trajectory $\bm{Z}$ (using the generated states prior to forcing) are then used to compute the modality-specific negative log-likelihoods for the observed multimodal time series,
$\mathcal{L}_{\text{DSR}}=-\sum_{t=1}^{T}(\log p_{\bm{\theta}}(\bm{{x}_{t}}|\bm{z}_{1:K,t})+\log p_{\bm{\theta}}(\bm{o}_{t}|\bm{z}_{1:K,t})+\log p_{\bm{\theta}}(\bm{c}_{t}|\bm{z}_{1:K,t})),$ (19)
using the decoder models from Eqs. 2-4.
Likewise, to connect the latent codes of the MVAE and the RNN, the model prior of the MVAE is instantiated through the RNN by taking $\bm{\mu}_{t}=\bm{z}_{1:K,t}$, i.e., only the first $K$ states of the generated latent sequence $\{\bm{z}_{t}\}$. As the initial state $\bm{z}_{1}$ is estimated directly from the encoded state $\bm{\tilde{z}}_{1}$, the term for $t=1$ evaluates to zero. Setting $L=1$, the consistency loss, eq. 9, between encoded and generated latent state paths thus comes down to
$\mathcal{L}_{\text{con}}=\frac{1}{2}\sum_{t=2}^{T}\bigl{(}\log|\bm{\Sigma}|+(\bm{\tilde{z}}_{t}-\bm{z}_{1:K,t})^{\top}\bm{\Sigma}^{-1}(\bm{\tilde{z}}_{t}-\bm{z}_{1:K,t})\bigl{)}.$ (20)
The total MTF loss is hence given by the loss in eq. 20 that ensures consistency between the latent codes of the MVAE and DSR model, the DSR model loss from the likelihoods of the observed time series $\bm{Y}$ given the predicted latent path $\bm{Z}$ (eq. 19), and the remaining terms from the ELBO, eq. 8 (noting that $\mathcal{L}_{\text{con}}=-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{\tilde{Z}})]$):
$\mathcal{L}_{\text{MTF}}=-\mathbb{E}_{q_{\bm{\phi}}}[\log p_{\bm{\theta}}(\bm{Y}|\bm{\tilde{Z}})]-\mathbb{H}_{q_{\bm{\phi}}}(\bm{\tilde{Z}}|\bm{Y})+\mathcal{L}_{\text{con}}+\mathcal{L}_{\text{DSR}}$ (21)
To train the dendPLRNN with MTF, RAdam (Liu et al., 2020) was used with a learning rate scheduler that iteratively reduced the learning rate from $10^{-3}$ to $10^{-5}$ during training. For each epoch, we randomly sampled sequences of length $T_{seq}=300$ from the total training data with a batch size of $16$, except for the fMRI data, where we chose $T_{seq}=72$ due to the short overall length of the data ($T=360$ per subject). The network weights $\bm{A}$, $\bm{W}$ and $\bm{h}$ from eq. 1 were initialized as suggested in Talathi & Vartak (2016). Expansion term thresholds $\{\bm{h}_{b}\}$ and input matrix entries $[\bm{U}]_{ij}$ were drawn from a standard normal distribution, while dendritic slopes $\alpha_{b}\sim\mathcal{U}(-B^{-\frac{1}{2}},B^{-\frac{1}{2}})$ were initialized uniformly.
A.2 Performance measures
Geometrical measure
To assess the (dis-)agreement $D_{\textrm{stsp}}$ between the data distribution $p_{\text{true}}({\bm{x}})$ and the model-generated distribution $p_{\text{gen}}({\bm{x}}|{\bm{z}})$ across state space, $\hat{p}_{\text{true}}({\bm{x}})$ and $\hat{p}_{\text{gen}}({\bm{x}}|{\bm{z}})$ are estimated by sampling $100$ trajectories with randomly drawn initial conditions and $1000$ time steps each. Transients are removed from each sampled trajectory to ensure that the (occupation) measure is evaluated on the limit set. The match between distributions is then approximated by binning the state space into discrete bins (Koppe et al., 2019b):
$D_{\mathrm{stsp}}\left(p_{\mathrm{true}}({\bm{x}}),p_{\mathrm{gen}}({\bm{x}}|{\bm{z}})\right)\approx\sum_{k=1}^{K}\hat{p}_{\mathrm{true}}^{(k)}({\bm{x}})\log\left(\frac{\hat{p}_{\mathrm{true}}^{(k)}({\bm{x}})}{\hat{p}_{\mathrm{gen}}^{(k)}({\bm{x}}|{\bm{z}})}\right).$ (22)
Here, $K$ is the total number of bins. A range of 2×2$\times$2 × the data standard deviation on each dimension was partitioned into $m$ bins, leading to a total of $K=m^{N}$ bins, where $N$ is the dimension of the ground truth system. Due to the exponential scaling of the number of bins with system dimensionality, for the $6$-dimensional Lewis-Glass network model and the fMRI data we instead used an approximation of $p_{\text{true}}({\bm{x}})$ and $p_{\text{gen}}({\bm{x}}|{\bm{z}})$ based on Gaussian mixture models placed along trajectories, as described in Brenner et al. (2022).
Geometric reconstruction measure in the absence of continuous observations
[…]
Power spectrum Hellinger distance
The power spectrum Hellinger distance ($D_{H}$) was obtained by first sampling a time series of $100,000$ time steps and computing dimension-wise Fast Fourier Transforms (using scipy.fft) for both the ground truth system and simulated time series, after discarding transients. The noise dominated high-frequency tails of the spectra were cut off, and the power spectra were slightly smoothed with a Gaussian kernel and normalized. We then computed the Hellinger distance (Mikhaeil et al., 2022) between smoothed power spectra of ground-truth, $F(\omega)$, and generated, $G(\omega)$, trajectories by
[…]
The dimension-wise Hellinger distances were then averaged to yield the $D_{H}$ values reported.
Spearman autocorrelation function (SACF)
To assess temporal agreement between model-generated and ground truth time series for ordinal and count observations, we computed a measure based on the average SACF. To this end we first sampled a time series of $100,000$ time steps and compute the dimension-wise Spearman autocorrelation for time lags up to $200$ for both generated and ground truth test data (see Fig. A9). The squared error between the resulting SACFs was then averaged across all dimensions.
Mean squared prediction error
A mean squared prediction error (PE) was computed across test sets of length $T=10000$ by initializing the trained dendPLRNN with the test set time series up to some time point $t$ from the estimated intial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$, from where it was then iterated forward by $n$ time steps to yield a prediction at time step $t+n$. The $n$-step PE is then defined as the MSE between predicted and true observations:
$PE(n)=\frac{1}{N(T-n)}\sum_{t=1}^{T-n}\sum_{i=1}^{N}(x_{i,t+n}-\hat{x}_{i,t+n})^{2}$ (24)
Due to exponential divergence of initially close trajectories in chaotic systems, the PE is sensible only for a limited number of time steps (Koppe et al., 2019b); here we chose $n=10$.
Determining initial states for computing mean squared prediction errors
Due to exponential divergence of initially close trajectories in chaotic systems, the prediction error (PE) also depends on the precise initialization at time $t$. For SVAE training, an estimate of the initial state is directly provided by the encoder model. For BPTT-TF, the initial state $\mathbf{z}_{t}$ can be inferred by a jointly trained linear mapping from $\mathbf{z}_{t}$ to $\mathbf{x}_{t}$ (Brenner et al., 2022). For MTF and GVAE-TF, if $K=M$, the initial state $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{1:t}]$ is also directly given by the encoder model. If $K<M$, a subset of $M-K$ latent states is drawn from a standard normal distribution. In this under-specified case, we used a warm-up phase of $t_{w}$ time steps, where the system was initialized from the encoder at time $t-t_{w}$ and iterated forward, providing encoded states in the form of TF signals at every time step, which yields an initial state estimate $\mathbf{E}[\mathbf{z}_{t}|\mathbf{x}_{(t-t_{w}):t}]$. We determined $t_{w}=20$ by grid-search, leading to on average best results across datasets. We found that including this warm-up phase significantly improved prediction performance, although it still led to slightly worse results than for the fully specified case, $K=M$. Finally, for MS we neither have an encoder model nor a TF signal, and hence a sensible estimate for the initial state is difficult to obtain.
Ordinal prediction error
The ordinal PE was computed similarly as the mean squared PE above, but – as pointed out in Öğretir et al. (2022) – due to the non-metric nature of ordinal data taking the absolute ($L_{1}$) deviation between observed and predicted values is more sensible:
$OPE(n)=\frac{1}{N(T-n)}\sum_{t=1}^{T-n}\sum_{i=1}^{N}|o_{i,t+n}-\hat{o}_{i,t+n}|$ (25)
Spearman cross-correlation (SCC)
To assess whether the global cross-correlation structure between the different ordinal time series is preserved by the reconstruction method, the Spearman correlation between each pair of ordinal time series was computed based on $100,000$ time steps long samples, using scipy.stats.spearmanr, for both generated and ground truth test data. The mean squared error between all elements of the correlation matrices was then taken.
Lyapunov exponent
The maximum Lyapunov exponent quantifies the divergence rate of nearby trajectories, and (for discrete time systems $F_{\bm{\theta}}\big{(}{\bm{z}}_{t-1},{\bm{s}}_{t}\big{)}$ like the dendPLRNN, Eq.1) is defined as
$\lambda_{max}:=\lim_{T\rightarrow\infty}\frac{1}{T}\log\left\lVert\ \prod_{r=0}^{T-2}{\bm{J}}_{T-r}\ \right\rVert_{2},$ (26)
where ${\bm{J}}_{t}:=\frac{\partial{\bm{F}}_{\bm{\theta}}\big{(}{\bm{z}}_{t-1},{\bm{s}}_{t}\big{)}}{\partial{\bm{z}}_{t-1}}=\frac{\partial{\bm{z}}_{t}}{\partial{\bm{z}}_{t-1}}$ are the system’s Jacobians and $\left\lVert\cdot\right\rVert_{2}$ is the spectral norm. To numerically approximate the Lyapunov spectrum for trained models, dendPLRNNs were iterated forward by $5500$ time steps using eq. 1, of which the first $500$ steps were discarded to remove transients. For numerical stability, an algorithm based on Wolf et al. (1985); Vogt et al. (2022) was used which re-orthogonalizes the product series of Jacobians after every $5$-th time step using a QR decomposition. For consistency, we also computed maximum Lyapunov exponents for the Lorenz ($\lambda_{max}=0.903$), Rössler ($\lambda_{max}=0.071$), and Lewis-Glass ($\lambda_{max}=0.072$) models ourselves using the Julia library DynamicalSystems.jl (Datseris, 2018) and the dysts Python package (Gilpin, 2022), which both use the same algorithm by Wolf et al. (1985), but our estimates agree closely with those that can be found in the literature (Lorenz: $\lambda_{max}=0.905$, Rössler: $\lambda_{max}=0.072$, in Alligood et al. (1996))
Figure A1: (a) Ground truth and rotated attractors of the Rössler system with associated $D_{stsp}$-values. (b) Correlation between geometrical reconstruction measures for the Rössler system directly in observation space given a co-trained linear (Gaussian) observation model ($D_{bin}$), and from a 3d PCA projection of latent space followed by an optimal rotation of the reconstructed attractor ($D_{PCA}$), based on a total of $30$ trained models.
A.3 Comparison methods
As noted further above, different choices of DSR model are possible within our framework (see Suppl. Table B3). Since the focus here is on evaluating the training framework for DSR from multimodal data itself, for comparability we used the same dendPLRNN model in all five other methods compared to. The total number of trainable parameters was kept approximately the same for all methods (see also Table B1).
Sequential MVAE
First, we compared our method to a sequential MVAE as proposed in (Kramer et al., 2022), optimizing its multimodal ELBO. This, to our knowledge, is currently the only other general approach specifically designed for DSR from arbitrary data modalities observed simultaneously. We followed the implementation of the encoder model as provided on https://github.com/DurstewitzLab/mmPLRNN, training with a sequence length of $150$ time steps per batch, a hidden dimension of $20$, and other model parameters similar to the runs with MTF.
Classical RNN training with truncated BPTT
Second, we tried ‘classical’ RNN training via truncated BPTT, where observations are provided as inputs at every time step, and the whole model is optimized using modality specific decoder models, eq. 19. Specifically, given observed multimodal time series $\bm{Y}$ of length $T$, observations $\bm{y}_{t}$ are included into the RNN updates according to
$\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\phi(\bm{z}_{t-1})+\bm{Uy}_{t}+\bm{h}+\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{\Sigma})$ (27)
where $\bm{U}$ is the input-to-hidden weight matrix. Training is performed on short subsequences of length $T_{seq}$, where $T_{seq}$ is a hyperparameter. The network then produces observations at each time step using the modality-specific decoder models, just as in MTF. During inference, the network is iterated forward by using the predicted observations at time $t$ as input for the predictions at time $t+1$. We probed different values for the sequence length $T_{seq}\in\{5\dots 50\}$. We found best settings to agree with the TF interval $\tau$ for MTF and the sequence length $L$ for multiple shooting, but – consistent with previous work (Mikhaeil et al., 2022) – plain error gradient/ sequence truncation often led to divergence and failed to capture the true time scales in the fMRI data.
Multiple shooting
A third approach that can deal with multimodal observation models but, unlike TF, does not require model inversion, is ‘multiple shooting (MS)’ (Bock & Plitt, 1984), a method suggested in the dynamical systems literature (Voss et al., 2004). Multiple shooting aims to solve boundary value problems by dividing time into sub-intervals, treating each sub-interval as a separate initial value problem, and then imposing continuity conditions between intervals. Multiple shooting has also been applied to DS reconstruction and Neural-ODE training (Iakovlev et al., 2022), where the initial conditions (‘shooting nodes’) for each interval are model parameters, and continuity across intervals is enforced through a penalty term in the loss (Voss et al., 2004). Hence, rather than controlling trajectory flows through a sparse TF signal applied after forcing intervals $\tau$, alternatively one may reset the latent model trajectory to an inferred initial condition after $\tau$ time steps. The advantage is that this method does not require inversion of decoder models and is hence naturally suited to handle different data modalities without further care (i.e., retaining the distributional properties of the original data). More specifically, the observed time series $\bm{Y}$ is partitioned into $N_{seq}$ subsequences $\bm{Y}^{s},s=1\dots N_{seq}$, of length $L$, and for each subsequence a new initial condition $\bm{\mu}_{0}^{s}$ is learned. During training, trajectories are freely generated for $L$ time steps from $\bm{\mu}_{0}^{s}$ for each subsequence, and likelihoods for the observed trajectories $\bm{Y}^{s}$ are computed using the observation models from Eqs. 2-4. A consistency (penalty) term in the loss ensures continuity between subsequences according to
$\mathcal{L}_{\text{MS}}=\lambda_{\text{MS}}\sum_{s=1}^{N_{seq}-1}||F_{\theta}(\bm{z}_{L}^{s})-\bm{\mu}_{0}^{s+1}||_{2}^{2}$ (28)
where $F_{\theta}$ in our case is the dendPLRNN, eq. 1, $\lambda_{\text{MS}}$ is a regularization parameter, and $F_{\theta}(\bm{z}^{s}_{L})=F_{\theta}(F_{\theta}(\dots F_{\theta}(\bm{\mu}_{0}^{s})))=F_{\theta}^{L}(\bm{\mu}_{0}^{s})$. The sequence length $L$ plays a similar role as the teacher forcing interval $\tau$ for MTF, controlling the times at which states and gradients are reset during training. Indeed, best performance is achieved when using choosing $L=\tau$ for all datasets studied here (see Table B1).
Standard unimodal approach with data ‘Gaussianization’
A naive approach for handling multimodal observations with any type of DSR model would be to pre-process all modalities such as to bring them into approximate agreement with Gaussian assumptions. Thus, for training the dendPLRNN with standard BPTT-TF (Brenner et al., 2022) and GVAE-TF, we transformed ordinal and count observations into approximately Gaussian variables through a Box-Cox-transformation (Box & Cox, 1964), $z$-scoring, and Gaussian kernel smoothing across the time series. For the optimal width of the Gaussian kernel, we performed a grid search over kernel sizes $\nu\in\{0,0.01,0.1,1,10,15,20,25\}$. Optimal settings for the results reported are given in Table B1.
A.4 Alternative encoder models
Mixture-of-Experts CNN encoder
The Mixture-of-Experts (MoE) approach (Shi et al., 2019) employs the same architecture as the CNN encoder described in Methods A.1, but uses a distinct encoder for each modality. We combined the outputs of each encoder into a joint estimate using a weighted average:
$\bm{\mu}_{\text{MoE}}$ $=w_{g}\bm{\mu}_{g}+w_{o}\bm{\mu}_{o}+w_{c}\bm{\mu}_{c}$
$\bm{\Sigma}_{\text{MoE}}$ $=w_{g}\bm{\Sigma}_{g}+w_{o}\bm{\Sigma}_{o}+w_{c}\bm{\Sigma}_{c},$
with means $\bm{\mu}_{\{g,o,c\}}$, diagonal covariances $\bm{\Sigma}_{\{g,o,c\}}$ and weights $w_{g}$ (Gaussian), $w_{o}$ (ordinal) and $w_{c}$ (count process), all set to $1/3$. We also tested a product-of-experts (PoE) approach (Hinton, 2002; Wu & Goodman, 2018), but found that it led to numerical instabilities for ordinal variables.
RNN encoder
For the RNN encoder (Cho et al., 2014), the hidden states $\bm{h}_{t}$ of an RNN, where observations $\bm{y}_{t}$ are provided as inputs at every time step (similar to eq. 27), are used to map onto the parameters of the approximate posterior at every time step. Here we use the standard RNN implementation in torch.nn.RNN, where
$\bm{h}_{t}$ $=\tanh(\bm{W}\bm{h}_{t-1}+\bm{U}\bm{y}_{t}+\bm{b})$
$\bm{\mu}_{t}$ $=\bm{W}_{\mu}\bm{h}_{t}+\bm{b}_{\mu}$
$diag\left([\log\sigma_{1}^{2},\ \dots,\ \log\sigma_{K}^{2}]\right)$ $=\bm{W}_{\Sigma}\bm{h}_{t}+\bm{b}_{\Sigma}$
with RNN parameters $\{\bm{W},\bm{U},\bm{b}\}$ and linear readout weights $\{\bm{W}_{\mu},\bm{b}_{\mu}\}$ for the mean and $\{\bm{W}_{\Sigma},\bm{b}_{\Sigma}\}$ for the logarithm of the diagonal covariance of the approximate posterior, respectively.
Transformer
The Transformer encoder is based on the architecture in (Vaswani et al., 2017). Since computation time scales with $T_{seq}^{2}$ in this architecture, we restricted the sequence length to $100$ steps. We used positional encodings to acknowledge the sequential (time series) nature of the data, as proposed in (Vaswani et al., 2017). The sequence plus embeddings is then passed through a standard Transformer encoder block using torch.nn.TransformerEncoder. As for the other encoder models, the output of the Transformer was mapped via linear readout layers onto the mean and logarithm of the covariance of the approximate posterior.
Multi-layer-perceptron (MLP)
We also tested a standard MLP with $3$ layers and rectified linear unit (ReLU) nonlinearity as encoder model, with outputs mapped via a linear readout layer to the mean and logarithm of the covariance of the approximate posterior, as for the other encoders tested.
Performance comparisons of these different encoder architectures are provided in Table B2.
A.5 Datasets
Lorenz-63 system
A stochastic version of the $3$d Lorenz-63 system, originally proposed in Lorenz (1963), is defined by
$\mathrm{d}x$ $=(\sigma(y-x))\mathrm{d}t+\mathrm{d}\epsilon_{1}(t),\,$
$\mathrm{d}y$ $=(x(\rho-z)-y)\mathrm{d}t+\mathrm{d}\epsilon_{2}(t),\,$ (29)
$\mathrm{d}z$ $=(xy-\beta z)\mathrm{d}t+\mathrm{d}\epsilon_{3}(t).$
Parameters used for producing ground truth data in the chaotic regime were $\sigma=10,\rho=28$, and $\beta=8/3$. Process noise was injected into the system by drawing from a Gaussian $\mathrm{d}\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\mathrm{d}t\times\bm{I})$. For both training and test data, a trajectory of $100,000$ time steps was sampled, performing numerical integration with scipy.odeint ($\mathrm{d}t=0.05$; note this value differs from the one used in Mikhaeil et al. (2022), explaining the different $\tau$ values required). To obtain multimodal observations, trajectories drawn from the ground truth system were fed into the different types of observation models in Eqs. 2-4, with randomly drawn parameters.
Symbolic representation of Lorenz-63 dynamics
To obtain a symbolic representation of the Lorenz-63 system, we first sampled a time series as above, using eq. A.5, and removed transients. We then divided the system’s state space into $N^{3}$ cubes (with $N=4$ bins per dimension). The range on each axis covered by the bins was set to the minimum-to-maximum extent of the attractor. Each data point of the time series was then assigned a label corresponding to the cube it fell in, using a one-hot encoding of length $N^{3}$. In fact, as for the Lorenz-63 chaotic attractor $28$ of the $64$ bins turned out empty (did not contain any points of the time series), the number of symbols (i.e., vector length) could be reduced further to just $36$ (see also Fig. 3 for decoded bin probabilities). The procedure is sketched in the bottom half of Fig. 3. We then furthermore performed a delay-embedding of the symbolic time series using a delay $\tau=10$ and an embedding dimension of $d=2$, thus doubling the length of the symbolic vector ($d=3$ gave similar results; see Matilla-García et al. (2021) for a discussion of how to determine optimal delay embedding parameters for symbolic sequences, although not used for the present purposes). Although not strictly necessary, this step profoundly improved DS reconstruction from symbolic sequences.
Rössler system
The Rössler system was introduced in Rössler (1976) as a simplified version of the Lorenz system, and is given (in SDE form) by
$\mathrm{d}x$ $=(-y-z)\mathrm{d}t+\mathrm{d}\epsilon_{1}(t),\,$
$\mathrm{d}y$ $=(x+ay)\mathrm{d}t+\mathrm{d}\epsilon_{2}(t),\,$ (30)
$\mathrm{d}z$ $=(b+z(x-c))\mathrm{d}t+\mathrm{d}\epsilon_{3}(t).$
Parameters used for producing ground truth data in the chaotic regime were $a=0.2,b=0.2$, and $c=5.7$. Process noise was added by drawing $\mathrm{d}\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\mathrm{d}t\times\bm{I})$. Training and test data was sampled as described above for the Lorenz-63 system, using $\mathrm{d}t=0.1$.
Lewis-Glass chaotic network model
We simulate a $6$-dimensional model of a neural network, originally introduced in Lewis & Glass (1992). Here the individual units of the network are endowed with a continuous gain function $G(x)=\frac{1+\tanh(-\alpha x)}{2}$, with the vector field given by
$\frac{\mathrm{d}\bm{x}}{\mathrm{d}t}=\frac{-\bm{x}}{\tau}+G(\epsilon\,\bm{K}\bm{x})-\beta.$ (31)
To sample from this system in the chaotic regime, we used the Hopfield model implementation in the Python package dysts.flows, based on Gilpin (2022). Here, $\alpha=-1,\beta=0.5,\epsilon=10,\tau=2.5$, and
$\bm{K}=\begin{bmatrix}0&-1&0&0&-1&-1&\\ 0&0&0&-1&-1&-1&\\ -1&-1&0&0&-1&0&\\ -1&-1&-1&0&0&0&\\ -1&-1&0&-1&0&0&\\ 0&-1&-1&-1&0&0&\\ \end{bmatrix}$
We generated training and test data by maketrajectory, and down-sampled the generated data by a factor of $30$. We sampled ordinal and count data in the same way as for the other datasets. Example time series and reconstructions are displayed in Fig. A12.
Human fMRI dataset
This same dataset was used previously in the study by Kramer et al. (2022) on multimodal data integration for DS reconstruction, and is openly available at https://github.com/DurstewitzLab/mmPLRNN. In the study, $26$ participants (of which $20$ were selected for analysis, cf. (Kramer et al., 2022)) were shown a series of images of different shapes (rectangles and triangles) while undergoing fMRI with a sampling rate of $1/3$ $\mathrm{Hz}$. The subjects were asked to identify the type of shape presented in the current image or in the preceding one, respectively, under three different task conditions: a continuous delayed response 1-back task (CDRT), a continuous matching 1-back task (CMT), and a 0-back control choice reaction task (CRT). A resting condition and an instruction phase were also included. Each task condition was repeated five times, where the last repetition of all task stages was left out as a test set in our analyses (see Fig. A2(b)). The brain activity of the participants as assessed via the BOLD signal, projected onto the first principal component within each of $10$ different brain regions in both hemispheres that were identified to be relevant for this the task. More details on this study can be found in Koppe et al. (2014). Due to the short length of these time series, for the fMRI results from Table 2 we chose a sequence length of $72$ time steps for testing and training.
Hippocampal multiple single-unit (MSU) and spatial position data
For our second empirical example, we used electrophysiological recordings from the rodent hippocampus and spatial location data (Grosmark et al., 2016), publicly available at https://crcns.org/data-sets/hc/hc-11/about-hc-11. Specifically, the “sessInfo.mat” file from rat ‘Achilles’, which contains extracted spike times, was chosen for exemplification and further preprocessed using the script from (Zhou & Wei, 2020), provided at https://github.com/zhd96/pi-vae/blob/main/code/rat_preprocess_data.py, to obtain counts per time interval, using a binning width of $200$ ms. For our analysis, we focused exclusively on the MAZE task and selected the $60$ most active neurons, as many neurons had very low (statistically insufficient) activity levels. The position data was represented as a continuous $1$d vector, with missing values imputed from neighboring points. Rats received water rewards at both end points of the track, which were provided to the model as scalar external signals $s_{t}$ (cf. eq. 1) set to $1$ for 5 time bins around the time points at which the rat started to move away from the reward location (with trainable weights $\bm{U}\in\mathbb{R}^{M\times 1}$ in eq. 1).
Appendix B Supplemental Results
Figure A2: (a) Multimodal integration using MTF on functional magnetic resonance imaging (fMRI)+behavioral data significantly helps to improve DS reconstruction compared to just fMRI data alone (unimodal). Results are for $20$ subjects (black lines; blue line = mean), showing geometrical ($D_{stsp}$, left) and temporal ($D_{H}$, right) disagreement between true and reconstructed systems; $p$-values from paired t-tests as indicated. (b) Example of decoded (color-coding of time series) and true (pale background colors) task stages for one brain area and subject. The trained model was freely iterated forward from the first time step of the test set, and task stages were decoded from the simulated activity. Decoding of class label $\hat{l}\in${Rest, Instruction, CRT, CDRT, CMT} is based on the maximum posterior probability, $\hat{l}_{t}={\arg\max}\ p(l_{kt}|\bm{z}_{t})$, given the latent trajectory $\bm{z}_{t}$. Graph below shows true and predicted BOLD signals. (c) 2d subspaces of freely generated latent activity for a dendPLRNN ($M=30,B=10,K=20,\tau=7$) trained jointly on continuous and categorical data by MTF for one example subject. The color coding corresponds to the task labels predicted according to the maximum posterior probability given the latent state as in (b) at each time step. The latent space appears to be structured according to the different task stages.
Figure A3: Impact of different loss terms in eq. 10 on DS reconstruction quality, indicating that all loss components are crucially important. Comparison of state space agreement $D_{stsp}$ (a) when excluding different terms from the total loss in eq. 10, and (b) for different scaling of the consistency loss ($\mathcal{L}_{\text{con}}$), for the dendPLRNN trained by MTF on data from the chaotic Lorenz system as in 3.1.
Figure A4: MTF loss landscapes (Eq. 10) as a function of two randomly chosen parameters ($A_{22}$, $W_{34}$) of the DSR model (Eq. 1) for four different settings of the TF interval $\tau$. The loss was computed on a random batch of training data for the model from Fig. 2. Smoother loss landscapes are obtained for lower $\tau$ (similar as observed for GTF in Hess et al. (2023)). While for the optimal TF interval $\tau=10$ the loss is smooth and convex, for too low $\tau=1$ flat directions in the loss appear.
Figure A5: Example reconstructions (best out of $10$ trained models each) of a dendPLRNN trained via MTF, compared to highly noisy Gaussian and ordinal training data (orange) produced by the Lorenz system. Reconstruction quality crucially depends on the choice of an optimal forcing interval, similar as observed for STF in (Mikhaeil et al., 2022). Choosing the interval too small (left) or too large (right) leads to significantly worse reconstructions.
Figure A6: Top: DS reconstruction from eight ordinal observations with seven levels each produced by the Lorenz-63 system. Example traces of ordinal time series (center), and example reconstruction by MTF (right) ($M=20,B=10,K=15,\tau=10$), preserving the butterfly-geometry of the Lorenz attractor. Bottom: DS reconstruction of the Rössler attractor from from $75$ ordinal observations with two levels each.
Figure A7: Agreement of true and predicted class label probabilities from a symbolic encoding of the reconstructed chaotic Lorenz-63 dynamics. Bins are color and transparency coded according to their respective relative frequencies. For the reconstructed system, category predictions correspond to those with the maximum posterior probability given a freely generated latent sequence (Eq. 15) sampled from a dendPLRNN trained with MTF ($M=30,B=15,K=30,\tau=10$).
Figure A8: Kernel-density estimates (using the default implementation of the kdeplot function in the seaborn library (Waskom, 2021)) of (a) distribution of maximum Lyapunov exponents, and (b) cumulative distributions of geometrical agreement $D_{stsp}$, for reconstructions produced by the MTF, Multiple Shooting (MS), and sequential multimodal VAE (SVAE), all trained with the same dendPLRNN model ($M=30,\ B=15,\ K=30$) on a symbolic encoding of the Lorenz-63 (eq. A.5). While for the MTF the estimates of Lyapunov exponents are centered around the true value ($\lambda_{max}=0.903$), for SVAE and MS the distributions were much farther off.
Figure A9: (a) Freely generated example trajectories and time series from a dendPLRNN ($M=20,\ B=10,\ K=15,\ \tau=10$) trained with MTF jointly on Gaussian, ordinal, and count data sampled from a Lorenz-63 system. (b) Example power spectra (Gaussian data) and Spearman autocorrelation functions (ordinal and count data). Simulated latent trajectories faithfully capture the geometry of the Lorenz attractor, as well as the temporal structure of the ground truth data when projected back into observation space.
Figure A10: Freely generated time series of a dendPLRNN ($M=30,B=15,K=30,\tau=7$) trained with MTF on continuous fMRI signals and categorical task variables from subjects a) #2, (b) #3, (c) #4, and (d) #7. Shown are time series from $10$ brain areas per subject. The trained dendPLRNN captures the overall temporal structure of the complex activity patterns from very short time series ($288$ time steps of training data for each subject).
Figure A11: Example reconstructions of spike trains and spike statistics on the training set (see Methods A.5). Test set reconstructions and further statistics are in Figure 4.
Figure A12: Example ground truth time series and freely generated series from a dendPLRNN ($M=20,B=10,K=15,\tau=20$) trained with MTF on the Lewis-Glass neural network model (Lewis & Glass, 1992).
Table B1: Hyperparameter settings for MTF, SVAE, BPTT, GVAE-TF, and MS trained on the Lorenz, Rössler and Lewis-Glass model, and for the experimental fMRI data. For explanation of symbols, see Methods A.3.
Dataset M B K $\tau,L,T_{seq}$ $\lambda_{\text{MS}}$ $\nu$
Lorenz $20$ $15$ $15$ $10$ $1.0$ $10$
Rössler $20$ $15$ $15$ $10$ $1.0$ $15$
Lewis-Glass $20$ $15$ $15$ $20$ $1.0$ $20$
fMRI $30$ $40$ $30$ $7$ $1.0$ $1$
Table B2: Performance comparison of different encoder models used in the MTF framework, trained on multimodal data from the chaotic Lorenz system. PE is the prediction error and OPE the ordinal prediction error. Note that SCC (Spearman cross-correlation), OACF (ordinal autocorrelation function), and CACF (count autocorrelation function) all refer to mean-squared-errors (MSEs) between ground truth and generated correlation functions.
Encoder $D_{stsp}$downarrow $D_{H}$downarrow PE $\downarrow$ OPE $\downarrow$ SCC $\downarrow$ OACF $\downarrow$ CACF $\downarrow$
CNN $\mathbf{3.4\pm 0.35}$ $\mathbf{0.30\pm 0.06}$ $\mathbf{1.3\mbox{e$-$}2}\pm\mathbf{2\mbox{e$-$}4}$ $\mathbf{0.12\pm 0.03}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{6.6\mbox{e$-$}5\pm 8\mbox{e$-$}6}$
CNN-MoE $5.89\pm 0.18$ $0.43\pm 0.03$ ${2.3\mbox{e$-$}2}\pm{5\mbox{e$-$}4}$ $0.13\pm 0.00$ $0.10\pm 0.00$ $0.19\pm 0.01$ ${1.1\mbox{e$-$}4\pm 2\mbox{e$-$}5}$
RNN $5.47\pm 0.48$ $0.32\pm 0.04$ ${1.6\mbox{e$-$}2}\pm{2\mbox{e$-$}4}$ $0.15\pm 0.01$ $0.13\pm 0.02$ $0.05\pm 0.01$ ${8.5\mbox{e$-$}5\pm 9\mbox{e$-$}6}$
Transformer $5.85\pm 0.14$ $0.40\pm 0.04$ ${4.8\mbox{e$-$}2}\pm{5\mbox{e$-$}4}$ $0.16\pm 0.00$ $0.17\pm 0.03$ $0.16\pm 0.02$ ${9.5\mbox{e$-$}5\pm 7\mbox{e$-$}6}$
MLP $6.57\pm 0.14$ $0.43\pm 0.01$ ${5.4\mbox{e$-$}2}\pm{6\mbox{e$-$}4}$ $0.15\pm 0.00$ $0.15\pm 0.01$ $0.21\pm 0.01$ ${1.3\mbox{e$-$}4\pm 9\mbox{e$-$}6}$
Table B3: Performance comparison of different RNNs used as DSR model in the MTF framework, incl. the dendPLRNN (Brenner et al., 2022) used here, an LSTM (Hochreiter & Schmidhuber, 1997), and a GRU (Chung et al., 2014), all trained by MTF on multimodal data from the chaotic Lorenz system ($10\%$ Gaussian observation noise). PE is the prediction error and OPE the ordinal prediction error. Note that SCC (Spearman cross-correlation), OACF (ordinal autocorrelation function), and CACF (count autocorrelation function) all refer to mean-squared-errors (MSEs) between ground truth and generated correlation functions.
RNN Model $D_{stsp}$downarrow $D_{H}$downarrow PE $\downarrow$ OPE $\downarrow$ SCC $\downarrow$ OACF $\downarrow$ CACF $\downarrow$
dendPLRNN $\mathbf{3.4\pm 0.35}$ $\mathbf{0.30\pm 0.06}$ $\mathbf{1.3\mbox{e$-$}2\pm 2\mbox{e$-$}4}$ $\mathbf{0.12\pm 0.03}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{0.07\pm 0.01}$ $\mathbf{6.6\mbox{e$-$}5\pm 8\mbox{e$-$}6}$
LSTM $3.8\pm 0.74$ $0.31\pm 0.01$ ${5.4\mbox{e$-$}2}\pm{5\mbox{e$-$}4}$ ${0.16\pm 0.03}$ $0.09\pm 0.02$ $0.09\pm 0.02$ ${8.8\mbox{e$-$}5\pm 8\mbox{e$-$}6}$
GRU $\mathbf{3.47\pm 0.56}$ $\mathbf{0.29\pm 0.03}$ ${3.5\mbox{e$-$}2}\pm{5\mbox{e$-$}4}$ ${0.13\pm 0.03}$ $\mathbf{0.06\pm 0.01}$ $\mathbf{0.08\pm 0.01}$ ${7.1\mbox{e$-$}5\pm 5\mbox{e$-$}6}$
Table B4: Comparison of dendPLRNN, eq. 1, trained by MTF (proposed method), by a sequential multimodal VAE (SVAE) based on (Kramer et al., 2022), and a multiple-shooting (MS) approach, on $8$ ordinal observations with seven ordered categories, produced by the chaotic Lorenz system, Rössler system, and Lewis-Glass model, and on a symbolic representation of the chaotic Lorenz system. Values are mean $\pm$ SEM, averaged over 15 trained models. X = value cannot easily be computed for MS (because here initial conditions cannot be obtained directly from the data but require additional parameters).
Dataset Method $D_{stsp}$downarrow $\lambda_{max}$ OPE $\downarrow$ SCC $\downarrow$ OACF $\downarrow$
Lorenz-ordinal MTF $\mathbf{8.8\pm 0.59}$ $\mathbf{0.92\pm 0.39}$ $\mathbf{0.24\pm 0.015}$ $\mathbf{0.085\pm 0.02}$ $\mathbf{0.016\pm 0.04}$
SVAE $14.7\pm 0.7$ $0.44\pm 0.71$ $0.8\pm 0.03$ $0.17\pm 0.02$ $0.23\pm 0.02$
MS $13.8\pm 1.1$ $0.47\pm 0.67$ X $0.24\pm 0.06$ $0.15\pm 0.03$
Rössler-ordinal MTF $\mathbf{7.9\pm 0.8}$ $\mathbf{0.03\pm 0.07}$ $\mathbf{0.093\pm 0.007}$ $\mathbf{0.051\pm 0.009}$ $\mathbf{0.051\pm 0.009}$
SVAE $11.5\pm 1.3$$0.27\pm 0.58$ $0.39\pm 0.02$ $0.23\pm 0.05$ $0.18\pm 0.04$
MS $14.1\pm 1.0$$0.05\pm 0.12$ X $0.12\pm 0.04$ $0.14\pm 0.03$
Lewis-Glass-ordinal MTF $\mathbf{0.89\pm 0.04}$ $\mathbf{-0.11\pm 0.41}$ $\mathbf{0.15\pm 0.02}$ $\mathbf{0.28\pm 0.05}$ $\mathbf{0.15\pm 0.03}$
SVAE $1.40\pm 0.22$ $-1.8\pm 2.1$ $0.29\pm 0.01$ $0.49\pm 0.04$ $0.24\pm 0.02$
MS $1.0\pm 0.14$ $-0.14\pm 0.31$ X $0.51\pm 0.04$ $0.45\pm 0.03$
Lorenz-symbolic MTF $\mathbf{4.4\pm 2.69}$ $\mathbf{0.67\pm 0.37}$
SVAE $12.02\pm 1.98$ $1.87\pm 0.88$
MS $\mathbf{4.46\pm 1.82}$ $5.67\pm 1.25$